OpenStudy (anonymous):

How do I find the domain of f(x) = (sqrt)(x^2-3x+2) ?

OpenStudy (across):

A function with a radical in it is not defined where the radicand is less than zero.

hero (hero):

x^2 - 3x + 2 = 0

OpenStudy (anonymous):

The equation is : \[f(x) = \sqrt{x^2-3x+2}\]

OpenStudy (across):

Instead of solving\[f(x)=\sqrt{x^2-3x+2}\]you solve\[x^2-3x+2\geq0\]That will give you the domain of your function.

hero (hero):

x^2 - 2x - 1x + 2 = 0 x(x-2)-1(x-2) = 0 (x-2)(x-1) =0 x = 2,1 x has to be greater than 2

OpenStudy (anonymous):

Do I factor to solve?

OpenStudy (across):

\[x^2-3x+2=(x-1)(x-2)\]

hero (hero):

has to be greater than or equal to 2

OpenStudy (anonymous):

So the domain would be [2, infinity) ?

OpenStudy (anonymous):

I think less than 1 works as well

hero (hero):

[-inf,1][2,inf]

OpenStudy (anonymous):

(- inf, 1] U [2, inf) ?

OpenStudy (anonymous):

exactly remidia

OpenStudy (across):

You're correct, remidia.

OpenStudy (anonymous):

thank you!

hero (hero):

I guess that implies that I was incorrect

OpenStudy (anonymous):

It does :.( You can't have [-inf, 1][2, inf] as you can never reach infinity

hero (hero):

Oh, I made a slight mistake there

OpenStudy (anonymous):

Can anyone help me with this? I need to find the domain of \[h(x) = 1/( sinx - (1/2))\]